OpenStudy (skullpatrol):

In the formula x/100 = a/total which is the percent: x or x/100?

OpenStudy (anonymous):

You say x% when it is x/100

OpenStudy (anonymous):

SO, when someone tells you x/100 It is x% Just x itself is not a percentage But when it is divided by 100 it becomes

OpenStudy (anonymous):

x/100 :)

OpenStudy (dan815):

yah per cent means per "100"

OpenStudy (skullpatrol):

But when asked to find what percent one number is of another, I answer x=(a/total)*100

OpenStudy (anonymous):

Take a example and ask yourself

OpenStudy (dan815):

^

OpenStudy (anonymous):

Variables are confusing

OpenStudy (dan815):

if u want to know what percent a is of b then = a/b * 100 =x% how much is 20 of 50 20/50*100 = 40%

OpenStudy (skullpatrol):

That is not the same as x/100 = a/b @dan815

OpenStudy (anonymous):

x/100 = a/total x means for 100 there are x so x/100 itsef is the percent (a/total )*100 =x is how much in 100 of sample ( for example )

OpenStudy (dan815):

when a/b is represented as a fraction with the denominator 100 then the numerator is the %

OpenStudy (skullpatrol):

The numerator is x @dan815

OpenStudy (anonymous):

so x define a value of material mmm its like aphysical thing lol it has a Module while x/100 is a percentage or ratio so there is uge diffrence btw them

hero (hero):

% = 1/100

OpenStudy (anonymous):

[x/100] is not percent here x is percent per - cent [ Per hundred] If [x/100] was the percent like [x/100]% then it would have become x/100/100 x/10000

hero (hero):

So therefore x% = x(1/100)

OpenStudy (skullpatrol):

To find what percent one number is of another, I answer $$\Huge x=\frac{a}{total} \cdot 100$$

OpenStudy (anonymous):

\[\Huge YES\]

hero (hero):

To further clarify, % = 1/100 = .01 therefore x% = x(.01) = .01x That's why went you take the percent of a number you have to move the decimal two places to the left.

OpenStudy (skullpatrol):

So x is a percent.

OpenStudy (anonymous):

if it is divided by 100 U can't take a random "x" and say it is percent

hero (hero):

x represents a number. % is 1/100 or .01 x% means a number multiplied by 1/100 hence x/100 It's as simple as that.

hero (hero):

So since x% = x/100 you can say that the expression x/100 means "x percent of a number"

OpenStudy (skullpatrol):

x/100 can not mean "x percent of a number"??? You haven't taken a percent of anything. @Hero

hero (hero):

I explained it as best I could. All I know is, I'm not the one confused. Just trying to help you out.

OpenStudy (skullpatrol):

Thanks @Hero :-)

hero (hero):

I probably should have said more clearly that x/100 simply means "percent of a number".

hero (hero):

and x represents some number of interest.

OpenStudy (skullpatrol):

I have heard a good interpretation that goes: percent is the numerator of a fraction whose denominator is 100 :-)